OpenStudy (anonymous):

intevsive math. 9th grade. 3(2x+1) for x= -8

OpenStudy (amistre64):

oh thats college algebra too

OpenStudy (anonymous):

in the USA

OpenStudy (amistre64):

and parts of venezuela maybe

OpenStudy (anonymous):

yea its supposed to CPM (collage prep i think)

OpenStudy (anonymous):

just sub the -8 in for x in the expression and evaluate

OpenStudy (riley):

Plug in -8 where you see x. 3(2(-8)+1) Then you work through it and solve everything out. 2 * -8 = -16 So your left with: 3(-16+1) -16+1 = -15 3(-15) = -45 So your answer is -45!

OpenStudy (anonymous):

okay so, ... 3(2x8+1) now what? I dont know what to do

OpenStudy (amistre64):

3(2x+1) for x= -8 3(2x+1) = N 2x+1 = N/3 x+1/2 = N/6 x= N/6 - 1/2 x= (N-3)/6 ; when x=-8, N-3 = -48 N = -45

OpenStudy (anonymous):

Thank you all for helping me!!!

OpenStudy (amistre64):

dios de machina lol

OpenStudy (riley):

Just make sure you always work the math out in the correct order of operations. PEMDAS = Please excuse my dear aunt sally. It's a funny saying to help you remember the order of operations. It stands for: "Parentheses, Exponents, Multiplication, Division, Addition, Subtraction"

OpenStudy (amistre64):

lol ... order of ops

OpenStudy (anonymous):

sub in the -8 for x\[3(2(-8)+1)\]simplify the mult inside the parens first\[=3(-16+1)\]continue to simplify inside the parens\[=3(-15)\]then multiply the 3\[=-45\]

OpenStudy (anonymous):

riley's post is VERY useful to know

OpenStudy (anonymous):

yea agreed. (:

OpenStudy (anonymous):

\[x-6 \over 4 - 1\]

OpenStudy (amistre64):

order of op it; and remember to divide before you add ro subtract ;)

OpenStudy (anonymous):

yea how do you apply that?